# A Bregman ADMM for Bethe variational problem

### v2

@@ 5.2 Quantum belief propagation
There are two types of quantum belief propagation (QBP) methods. The approach introduced by Leifer and Poulin in [35] is a direct extension of the classical belief propagation. However, it imposes stringent requirements on the quantum system and may not converge, even when the graph $G$ is a tree. In this paper, we focus on the QBP method proposed by Zhao et al. in [88], which is derived from the KKT conditions of (5.1) and is applicable to a broader class of problems.
@@
[…]
@@
^$\widehat{\mathcal{H}}_{1}:=\displaystyle\prod\limits_{k\in[n]}\mathbb{H}^{r},\ \widehat{\mathcal{H}}_{2}:=\displaystyle\prod\limits_{ij\in\mathcal{E}}\mathbb{H}^{r^{2}},\ \widehat{\mathcal{H}}_{3}:=\displaystyle\prod\limits_{ij\in\mathcal{E}}\mathbb{H}^{r}\times\mathbb{H}^{r}.$ (170)
@@
$\widehat{\mathcal{X}}$ ::=:= $\left\{{\bf q}\in\widehat{\mathcal{H}}_{1}:\ \forall k\in[n],\ q_{k}\in\mathbb{H}^{r}_{+},\ \mathrm{Tr}(q_{k})=1\right\},$ (171)
$\widehat{\mathcal{Y}}$ ::=:= $\left\{{\bf Q}\in\widehat{\mathcal{H}}_{2}:\ \forall ij\in\mathcal{E},\ Q_{ij}\in\mathbb{H}^{r^{2}}_{+},\ \mathrm{Tr}(Q_{ij})=1\right\},$ (172)
and linear maps $\widehat{\mathcal{A}}:\widehat{\mathcal{H}}_{1}\rightarrow\widehat{\mathcal{H}}_{3},$:^$\widehat{\mathcal{B}}:\widehat{\mathcal{H}}_{2}\rightarrow\widehat{\mathcal{H}}_{3}$ such that for any ${\bf q}\in\widehat{\mathcal{H}}_{1},$ ${\bf Q}\in\widehat{\mathcal{H}}_{2}$
@@
Define the functions ^$\widehat{\mathcal{F}}:\mathcal{X}\rightarrow\mathbb{R}$ and $\widehat{\mathcal{G}}:\mathcal{Y}\rightarrow\mathbb{R}$ such that for any ${\bf q}\in\widehat{\mathcal{X}},$ ${\bf Q}\in\widehat{\mathcal{Y}}$,
^$\widehat{\mathcal{F}}({\bf q}):=\left\langle\hat{{\bf c}},{\bf q}\right\rangle-\sum_{k\in[n]}(d_{k}-1)\left\langle q_{k},{\rm logm}(q_{k})\right\rangle,\ \widehat{\mathcal{G}}({\bf Q}):=\left\langle\widehat{{\bf C}},{\bf Q}\right\rangle+\sum_{ij\in\mathcal{E}}\left\langle Q_{ij},{\rm logm}(Q_{ij})\right\rangle.$ (174)
@@
Similar to the strong subadditivity (SSA) of Shannon entropy, von Neumann entropy also has SSA [41]. Thus, the following function $\hat{\varphi}:\widehat{\mathcal{Y}}\rightarrow\mathbb{R}$ is convex
@@ 6 Numerical experiments
For CCCP, as recommended in [87], we run the BCM subroutine for 5 iterations to solve its subproblem approximately. BP has two variants: Jacobi Belief Propagation (JBP) and Gauss-Seidel Belief Propagation (GSBP). JBP updates all messages simultaneously in parallel during each iteration, akin to the Jacobi method in numerical linear algebra. In contrast, GSBP updates each message sequentially, incorporating the most recent information from previously updated messages within the same iteration. Similarly, QBP has two variants JQBP and GSQBP.
@@ 6.1 Example of (BVP): spin-glass models
In the first example, we test spin-glass models, which has been tested in [87, 83, 78]. We generate problem data in the same way as in [87]. In detail, the graph $G$ is generated as an $n_{1}\times n_{1}$ 2D grid graph, where $n_{1}\in\{50,100\}$ and $n_{1}\times n_{1}\times n_{1}$ 3D lattice graph, where $n_{1}\in\{20,50\}.$ The data $c_{k}\in\mathbb{R}^{2}$ and $C_{ij}\in\mathbb{R}^{2\times 2}$ are generated from Gaussian distribution $N(0,\sigma^{2})$ where $\sigma\in\{1,2,5\}.$ The results are shown in Table LABEL:spin2d and Table LABEL:spin3d.
@@
$n$=2500 JBP* 3.58e+00* 0 -1.5038001e+04 10000 2.75e+01
$m$=4900 GSBP* 5.26e+00* 0 -1.5035182e+04 10000 1.70e+03
$\sigma$=5 BADMM 2.54e-07 7.19e-07 -1.5034679e+04 261 1.08e+00
@@
$n$=10000 JBP 4.80e-07 0 -1.3851321e+04 19 2.03e-01
$m$=19800 GSBP 2.73e-07 0 -1.3851320e+04 12 8.72e+00
$\sigma$=1 BADMM 2.90e-07 8.19e-07 -1.3851332e+04 171 2.02e+00
@@
$n$=10000 JBP 6.25e-07 0 -2.5208488e+04 67 6.57e-01
$m$=19800 GSBP 9.63e-07 0 -2.5208485e+04 31 2.24e+01
$\sigma$=2 BADMM 2.35e-07 8.27e-07 -2.5208490e+04 411 4.76e+00
@@
$n$=10000 JBP* 2.28e+01* 0 -6.1346285e+04 10000 9.70e+01
$m$=19800 GSBP* 4.02e+01* 0 -6.1344648e+04 5030 3.60e+03
$\sigma$=5 BADMM 3.44e-07 9.37e-07 -6.1351748e+04 401 4.70e+00
@@
$n$=8000 JBP 7.90e-07 0 -1.2344487e+04 26 3.35e-01
$m$=22800 GSBP 8.26e-07 0 -1.2344486e+04 15 1.35e+01
$\sigma$= 1 BADMM 4.31e-07 8.27e-07 -1.2344498e+04 321 3.81e+00
@@
$n$=8000 JBP* 4.69e-01* 0 -2.3002998e+04 10000 1.10e+02
$m$=22800 GSBP* 8.58e-01* 0 -2.3003425e+04 4237 3.60e+03
$\sigma$= 2 BADMM 5.25e-07 9.98e-07 -2.3003083e+04 471 5.54e+00
@@
$n$=8000 JBP* 9.47e+01* 0 -5.6343253e+04 10000 1.09e+02
$m$=22800 GSBP* 9.28e+01* 0 -5.6437373e+04 4273 3.60e+03
$\sigma$= 5 BADMM 5.03e-07 9.07e-07 -5.6441554e+04 711 8.02e+00
@@ 6.2 Example of (BVP): sensor network localization
In the second example, we test the Sensor Network Localization (SNL) problem [48], which aims to estimate the locations of sensors based on noisy observations of pairwise distances. Traditional approaches treat SNL as a nonlinear regression problem and employ algorithms such as gradient descent [71, 39] and semidefinite programming [9, 14] to solve it. In [27], Ihler et al. reformulated SNL as a graphical model and proposed the Nonparametric Belief Propagation (NBP) method as a solver.
Specifically, consider $n$ sensors $x_{1},x_{2},\ldots,x_{n}$ located in a planar region $[0,1]^{2}$. For any distinct $i,j\in[n]$, the noisy distance between sensors is observed with probability $P_{o}(x_{i},x_{j})$, given by:
@@
Unlike previous graphical models with finite states, the potential functions in (183) and (184) are continuous functions. To apply our algorithm, we discretize $[0,1]^{2}$ into a grid $\{0,1/t,2/t,\ldots,(t-1)/t,1\}^{2}$ and restrict the domain of the potential functions to this discrete set. The resulting number of states is $r=(t+1)^{2}$. Note that this discretization introduces significant model error if $t$ is too small and dramatically increases problem size if $t$ is too large. Ihler et al. [27] addressed this issue by using NBP with stochastic approximation to solve the continuous model directly, avoiding discretization. Here, we use discretization solely to test the behavior of our algorithm. Extending Bregman ADMM to the continuous setting is an interesting direction for future research.
Akin to the experiments in [27], we test SNL problems with and without outlier noise. The first experiment doesn’t contain outlier noise. We set $n=100$, $a=4$, $t=10$, $\sigma\in\{0.02,0.05,0.1\}$, and $R\in\{0.1,0.2\}$. All sensors and anchors are uniformly distributed in $[0,1]^{2}$. We use the rng(’default’) command in Matlab to ensure reproducibility of results. Since the discretization already introduces model error, it is unnecessary to solve the problems to high accuracy. We choose the tolerance to be $10^{-4}$ for all the methods. The results are shown in Table LABEL:SNL.
@@
Figure 1: Results for SNL without outlier noise. $\sigma=0.02,R=0.2,m=100.$ +’s are predicted sensors, ’s are real sensors, ✩’s are anchors. RMSD$=\sqrt{\sum_{k\in[n]}\|x_{k}-\hat{x}_{k}\|^{2}/n},$ where $\{\hat{x}_{k}:\ k\in[n]\}$ are predicted locations of sensors.
@@
Figure 2: Results for SNL with $5\%$ outlier noise. $\sigma=0.005,R=0.2,m=100.$ +’s are predicted sensors, ’s are real sensors, ✩’s are anchors. RMSD$=\sqrt{\sum_{k\in[n]}\|x_{k}-\hat{x}_{k}\|^{2}/n},$ where $\{\hat{x}_{k}:\ k\in[n]\}$ are predicted locations of sensors.
@@
In the third example, we test the 2D Ising models in (5.1), which has been tested in [88]. We generate $G$ as an $n_{1}\times n_{1}$ 2D grid graph. The coefficient matrices are generated as
$\forall k\in[n],\ \hat{c}_{k}:=\frac{1}{T}\left(h_{x}\begin{bmatrix}0&1\\ 1&0\end{bmatrix}-h_{z}\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}\right),$
@@
where $h_{x},h_{z},J,T$ are parameters. We choose $n_{1}\in\{10,30\}$ and $[h_{x},h_{z},J]=[1.05,0.5,1]$ or $[2.5,0,-1]$ as suggested in [88].
@@
$n$=100 QJBP 2.60e-07 0 -2.4386471e+02 6 3.92e-01
$m$=180 QGSBP 7.08e-07 0 -2.4386511e+02 24 1.28e+00
$T$=1 QBADMM 2.54e-07 4.15e-07 -2.4386494e+02 51 1.73e+00
$n$=100 QJBP* - - - - -
$m$=180 QGSBP 9.88e-07 0 -2.4416552e+03 436 2.11e+01
$T$=0.1 QBADMM 8.53e-07 2.97e-07 -2.4416553e+03 171 5.31e+00
$n$=100 QJBP* - - - - -
$m$=180 QGSBP 9.99e-07 0 -2.4478861e+04 3194 1.53e+02
$T$=0.01 QBADMM 9.71e-07 2.19e-07 -2.4478820e+04 591 1.81e+01
$n$=900 QJBP 4.50e-07 0 -2.3050728e+03 6 2.31e+00
$m$=1740 QGSBP 3.63e-07 0 -2.3050734e+03 25 1.09e+01
$T$=1 QBADMM 1.27e-07 1.96e-07 -2.3050722e+03 61 1.80e+01
$n$=900 QJBP* - - - - -
$m$=1740 QGSBP 9.99e-07 0 -2.3076988e+04 2415 1.16e+03
$T$=0.1 QBADMM 9.28e-07 2.80e-07 -2.3076992e+04 661 1.99e+02
$n$=900 QJBP* - - - - -
$m$=1740 QGSBP* 1.21e-04* 0 -2.3130566e+05 7355 3.60e+03
$T$=0.01 QBADMM 9.84e-07 2.31e-07 -2.3130470e+05 2211 6.62e+02
@@
$n$=100 QJBP* 5.81e+01* 0 -1.0040858e+02 10000 4.12e+02
$m$=180 QGSBP 7.06e-07 0 -2.7606344e+02 32 1.54e+00
$T$=1 QBADMM 1.01e-08 3.36e-08 -2.7606705e+02 51 1.70e+00
$n$=100 QJBP 3.50e-07 0 -2.5000487e+03* 13 4.80e-01
$m$=180 QGSBP 9.93e-07 0 -3.0046956e+03 902 4.25e+01
$T$=0.1 QBADMM 7.54e-07 2.58e-07 -3.0046879e+03 341 1.10e+01
$n$=100 QJBP 1.72e-07 0 -2.5000342e+04* 13 4.68e-01
$m$=180 QGSBP* 1.68e-05* 0 -3.0726971e+04 10000 4.63e+02
$T$=0.01 QBADMM 9.78e-07 2.05e-07 -3.0725593e+04 2741 8.48e+01
$n$=900 JQBP* 5.14e+02* 0 -5.2905916e+02 9295 3.60e+03
$m$=1740 QGSBP 6.18e-07 0 -2.5067091e+03 35 1.57e+01
$T$=1 QBADMM 1.33e-07 4.38e-07 -2.5067207e+03 51 1.49e+01
$n$=900 QJBP 8.25e-07 0 -2.2500146e+04* 13 4.19e+00
$m$=1740 QGSBP 9.99e-07 0 -2.7634677e+04 3083 1.36e+03
$T$=0.1 QBADMM 9.70e-07 2.14e-07 -2.7634590e+04 $1$ 3.25e+02
$n$=900 QJBP 4.02e-07 0 -2.2500103e+05* 13 4.20e+00
$m$=1740 QGSBP* 1.47e-03* 0 -2.8361552e+05 7824 3.60e+03
$T$=0.01 QBADMM* 2.55e-06* 5.25e-07 -2.8357454e+05 10000 2.92e+03
From Table LABEL:Ising_model_1, we observe that QBADMM is the only algorithm capable of solving all the instances. QJBP encounters numerical issues, and its output becomes NaN in Matlab, so we do not include its results in the table. This numerical instability has also been reported in the experiments conducted by Zhao et al. in [88]. Interestingly, QGSBP does not exhibit such numerical issues and is often able to find solutions with acceptable accuracy.

### v3

@@ 5.2 Quantum belief propagation
There are two types of quantum belief propagation (QBP) methods. The approach introduced by Leifer and Poulin in [leifer2008quantum] is a direct extension of the classical belief propagation. However, it imposes stringent requirements on the quantum system and may not converge, even when the graph $G$ is a tree. In this paper, we focus on the QBP method proposed by Zhao et al. in [zhao2024quantum], which is derived from the KKT conditions of (5.1) and is applicable to a broader class of problems.
@@
Here, $\prod\limits$ represents the sequence of $\odot$-products. After each round of message updates, the marginal density operator $q_{k}$ for any $k\in[n]$ is recovered as follows:
@@
$\widehat{\mathcal{H}}_{1}:=\displaystyle\prod\limits_{k\in[n]}\mathbb{H}^{r},\ \widehat{\mathcal{H}}_{2}:=\displaystyle\prod\limits_{ij\in\mathcal{E}}\mathbb{H}^{r^{2}},\ \widehat{\mathcal{H}}_{3}:=\displaystyle\prod\limits_{ij\in\mathcal{E}}\mathbb{H}^{r}\times\mathbb{H}^{r}.$ (170)
@@
$\widehat{\mathcal{X}}$ $:=$ $\left\{{\bf q}\in\widehat{\mathcal{H}}_{1}:\ \forall k\in[n],\ q_{k}\in\mathbb{H}^{r}_{+},\ \mathrm{Tr}(q_{k})=1\right\},$ (171)
$\widehat{\mathcal{Y}}$ $:=$ $\left\{{\bf Q}\in\widehat{\mathcal{H}}_{2}:\ \forall ij\in\mathcal{E},\ Q_{ij}\in\mathbb{H}^{r^{2}}_{+},\ \mathrm{Tr}(Q_{ij})=1\right\},$ (172)
and linear maps $\widehat{\mathcal{A}}:\widehat{\mathcal{H}}_{1}\rightarrow\widehat{\mathcal{H}}_{3},$ $\widehat{\mathcal{B}}:\widehat{\mathcal{H}}_{2}\rightarrow\widehat{\mathcal{H}}_{3}$ such that for any ${\bf q}\in\widehat{\mathcal{H}}_{1},$ ${\bf Q}\in\widehat{\mathcal{H}}_{2}$
@@
Define the functions $\widehat{\mathcal{F}}:\mathcal{X}\rightarrow\mathbb{R}$ and $\widehat{\mathcal{G}}:\mathcal{Y}\rightarrow\mathbb{R}$ such that for any ${\bf q}\in\widehat{\mathcal{X}}$, $^{\bf Q}\in\widehat{\mathcal{Y}}$,
$\widehat{\mathcal{F}}({\bf q}):=\left\langle\hat{{\bf c}},{\bf q}\right\rangle-\sum_{k\in[n]}(d_{k}-1)\left\langle q_{k},{\rm logm}(q_{k})\right\rangle,\ \widehat{\mathcal{G}}({\bf Q}):=\left\langle\widehat{{\bf C}},{\bf Q}\right\rangle+\sum_{ij\in\mathcal{E}}\left\langle Q_{ij},{\rm logm}(Q_{ij})\right\rangle.$ (174)
@@
Similar to the strong subadditivity (SSA) of Shannon entropy, von Neumann entropy also has SSA [lieb1973proof]. Thus, the following function $\hat{\varphi}:\widehat{\mathcal{Y}}\rightarrow\mathbb{R}$ is convex
@@ 6 Numerical experiments
For CCCP, as recommended in [yuille2002cccp], we run the BCM subroutine for 5 iterations to solve its subproblem approximately. BP has two variants: Jacobi Belief Propagation (JBP) and Gauss-Seidel Belief Propagation (GSBP). JBP updates all messages simultaneously in parallel during each iteration, akin to the Jacobi method in numerical linear algebra. In contrast, GSBP updates each message sequentially, incorporating the most recent information from previously updated messages within the same iteration. Similarly, QBP has two variants JQBP and GSQBP.
@@ 6.1 Example of (BVP): spin-glass models
In the first example, we test spin-glass models, which has been tested in [yuille2002cccp, yedidia2001bethe, welling2013belief]. We generate problem data in the same way as in [yuille2002cccp]. In detail, the graph $G$ is generated as an $n_{1}\times n_{1}$ 2D grid graph, where $n_{1}\in\{50,100\}$ and $n_{1}\times n_{1}\times n_{1}$ 3D lattice graph, where $n_{1}\in\{20,50\}.$ The data $c_{k}\in\mathbb{R}^{2}$ and $C_{ij}\in\mathbb{R}^{2\times 2}$ are generated from Gaussian distribution $N(0,\sigma^{2})$ where $\sigma\in\{1,2,5\}.$ The results are shown in Table LABEL:spin2d and Table LABEL:spin3d.
@@
nn=2500 JBP* 3.58e+00* 0 -1.5038001e+04 10000 2.75e+01
mm=4900 GSBP* 5.26e+00* 0 -1.5035182e+04 10000 1.70e+03
$\sigma$=5 BADMM 2.54e-07 7.19e-07 -1.5034679e+04 261 1.08e+$0$
@@
nn=10000 JBP 4.80e-07 0 -1.3851321e+04 19 2.03e-01
mm=19800 GSBP 2.73e-07 0 -1.3851320e+04 12 8.72e+$0$
$\sigma$=1 BADMM 2.90e-07 8.19e-07 -1.3851332e+04 171 2.02e+$0$
@@
nn=10000 JBP 6.25e-07 0 -2.5208488e+04 67 6.57e-01
mm=19800 GSBP 9.63e-07 0 -2.5208485e+04 31 2.24e+01
$\sigma$=2 BADMM 2.35e-07 8.27e-07 -2.5208490e+04 411 4.76e+$0$
@@
nn=10000 JBP* 2.28e+01* 0 -6.1346285e+04 10000 9.70e+01
mm=19800 GSBP* 4.02e+01* 0 -6.1344648e+04 5030 3.60e+03
$\sigma$=5 BADMM 3.44e-07 9.37e-07 -6.1351748e+04 401 4.70e+$0$
@@
nn=8000 JBP 7.90e-07 0 -1.2344487e+04 26 3.35e-01
mm=22800 GSBP 8.26e-07 0 -1.2344486e+04 15 1.35e+01
$\sigma$= 1 BADMM 4.31e-07 8.27e-07 -1.2344498e+04 321 3.81e+$0$
@@
nn=8000 JBP* 4.69e-01* 0 -2.3002998e+04 10000 1.10e+02
mm=22800 GSBP* 8.58e-01* 0 -2.3003425e+04 4237 3.60e+03
$\sigma$= 2 BADMM 5.25e-07 9.98e-07 -2.3003083e+04 471 5.54e+$0$
@@
nn=8000 JBP* 9.47e+01* 0 -5.6343253e+04 10000 1.09e+02
mm=22800 GSBP* 9.28e+01* 0 -5.6437373e+04 4273 3.60e+03
$\sigma$= 5 BADMM 5.03e-07 9.07e-07 -5.6441554e+04 711 8.02e+$0$
@@ 6.2 Example of (BVP): sensor network localization
In the second example, we test the Sensor Network Localization (SNL) problem [mao2007wireless], which aims to estimate the locations of sensors based on noisy observations of pairwise distances. Traditional approaches treat SNL as a nonlinear regression problem and employ algorithms such as gradient descent [tang2024riemannian, liang2004gradient] and semidefinite programming [biswas2004semidefinite, chaudhury2015large] to solve it. In [ihler2004nonparametric], Ihler et al. reformulated SNL as a graphical model and proposed the Nonparametric Belief Propagation (NBP) method as a solver.
Specifically, consider $n$ sensors $x_{1},x_{2},\ldots,x_{n}$ located in a planar region $0,1]^{2}$. For any distinct $i,j\in[n]$, the noisy distance between sensors is observed with probability $P_{o}(x_{i},x_{j})$, given by:
@@
Unlike previous graphical models with finite states, the potential functions in (183) and (184) are continuous functions. To apply our algorithm, we discretize $0,1]^{2}$ into a grid $\{0,1/t,2/t,\ldots,(t-1)/t,1\}^{2}$ and restrict the domain of the potential functions to this discrete set. The resulting number of states is $r=(t+1)^{2}$. Note that this discretization introduces significant model error if $t$ is too small and dramatically increases problem size if $t$ is too large. Ihler et al. [ihler2004nonparametric] addressed this issue by using NBP with stochastic approximation to solve the continuous model directly, avoiding discretization. Here, we use discretization solely to test the behavior of our algorithm. Extending Bregman ADMM to the continuous setting is an interesting direction for future research.
Akin to the experiments in [ihler2004nonparametric], we test SNL problems with and without outlier noise. The first experiment doesn’t contain outlier noise. We set $n=100$, $a=4$, $t=10$, $\sigma\in\{0.02,0.05,0.1\}$, and $R\in\{0.1,0.2\}$. All sensors and anchors are uniformly distributed in $0,1]^{2}$. We use the rng(’default’) command in Matlab to ensure reproducibility of results. Since the discretization already introduces model error, it is unnecessary to solve the problems to high accuracy. We choose the tolerance to be $10^{-4}$ for all the methods. The results are shown in Table LABEL:SNL.
@@
[…]
@@
[…]
@@
In the third example, we test the 2D Ising models in (5.1), which has been tested in [zhao2024quantum]. We generate $G$ as an $n_{1}\times n_{1}$ 2D grid graph. The coefficient matrices are generated as
[…]
@@
where $h_{x},h_{z},J,T$ are parameters. We choose $n_{1}\in\{10,30\}$ and $h_{x},h_{z},J]=[1.05,0.5,1]$ or [2.5,0,1][2.5,0,-1] as suggested in [zhao2024quantum].
@@
nn=100 QJBP 2.60e-07 0 -2.4386471e+02 6 3.92e-01
mm=180 QGSBP 7.08e-07 0 -2.4386511e+02 24 1.28e+$0$
TT=1 QBADMM 2.54e-07 4.15e-07 -2.4386494e+02 51 1.73e+$0$
nn=100 QJBP* - - - - -
mm=180 QGSBP 9.88e-07 0 -2.4416552e+03 436 2.11e+01
TT=0.1 QBADMM 8.53e-07 2.97e-07 -2.4416553e+03 171 5.31e+$0$
nn=100 QJBP* - - - - -
mm=180 QGSBP 9.99e-07 0 -2.4478861e+04 3194 1.53e+02
TT=0.01 QBADMM 9.71e-07 2.19e-07 -2.4478820e+04 591 1.81e+01
nn=900 QJBP 4.50e-07 0 -2.3050728e+03 6 2.31e+$0$
mm=1740 QGSBP 3.63e-07 0 -2.3050734e+03 25 1.09e+01
TT=1 QBADMM 1.27e-07 1.96e-07 -2.3050722e+03 61 1.80e+01
nn=900 QJBP* - - - - -
mm=1740 QGSBP 9.99e-07 0 -2.3076988e+04 2415 1.16e+03
TT=0.1 QBADMM 9.28e-07 2.80e-07 -2.3076992e+04 661 1.99e+02
nn=900 QJBP* - - - - -
mm=1740 QGSBP* 1.21e-04* 0 -2.3130566e+05 7355 3.60e+03
TT=0.01 QBADMM 9.84e-07 2.31e-07 -2.3130470e+05 2211 6.62e+02
@@
nn=100 QJBP* 5.81e+01* 0 -1.0040858e+02 10000 4.12e+02
mm=180 QGSBP 7.06e-07 0 -2.7606344e+02 32 1.54e+$0$
TT=1 QBADMM 1.01e-08 3.36e-08 -2.7606705e+02 51 1.70e+$0$
nn=100 QJBP 3.50e-07 0 -2.5000487e+03* 13 4.80e-01
mm=180 QGSBP 9.93e-07 0 -3.0046956e+03 902 4.25e+01
TT=0.1 QBADMM 7.54e-07 2.58e-07 -3.0046879e+03 341 1.10e+01
nn=100 QJBP 1.72e-07 0 -2.5000342e+04* 13 4.68e-01
mm=180 QGSBP* 1.68e-05* 0 -3.0726971e+04 10000 4.63e+02
TT=0.01 QBADMM 9.78e-07 2.05e-07 -3.0725593e+04 2741 8.48e+01
nn=900 JQBP* 5.14e+02* 0 -5.2905916e+02 9295 3.60e+03
mm=1740 QGSBP 6.18e-07 0 -2.5067091e+03 35 1.57e+01
TT=1 QBADMM 1.33e-07 4.38e-07 -2.5067207e+03 51 1.49e+01
nn=900 QJBP 8.25e-07 0 -2.2500146e+04* 13 4.19e+$0$
mm=1740 QGSBP 9.99e-07 0 -2.7634677e+04 3083 1.36e+03
TT=0.1 QBADMM 9.70e-07 2.14e-07 -2.7634590e+04 $11$ 3.25e+02
nn=900 QJBP 4.02e-07 0 -2.2500103e+05* 13 4.20e+$0$
mm=1740 QGSBP* 1.47e-03* 0 -2.8361552e+05 7824 3.60e+03
TT=0.01 QBADMM* 2.55e-06* 5.25e-07 -2.8357454e+05 10000 2.92e+03
From Table LABEL:Ising_model_1, we observe that QBADMM is the only algorithm capable of solving all the instances. QJBP encounters numerical issues, and its output becomes NaN in Matlab, so we do not include its results in the table. This numerical instability has also been reported in the experiments conducted by Zhao et al. in [zhao2024quantum]. Interestingly, QGSBP does not exhibit such numerical issues and is often able to find solutions with acceptable accuracy.